# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 2.1. Set up and logarithmic derivatives
i The logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ is an $\mathbb{R}^{d}$-valued function such that $\mathsf{d}^{\nu}\in L_{\mathrm{loc}}^{1}(\nu^{[1]})$ and that, for all $f\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}:=C_{0}^{\infty}(\mathbb{R}^{d})\otimes\mathscr{D}_{\circ\mathrm{b}}$,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}f(x,\mathsf{s})\mathsf{d}^{\nu}(x,\mathsf{s})\nu^{[1]}(dxd\mathsf{s})=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\frac{\partial}{\partial x}f(x,\mathsf{s})\nu^{[1]}(dxd\mathsf{s}).$ (2.10)
ii The $\bullet$-logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ is the $\mathbb{R}^{d}$-valued function defined similarly to the logarithmic derivative, but with $\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ replaced by $\mathscr{D}_{\bullet\mathrm{b}}^{[1]}:=C_{0}^{\infty}(\mathbb{R}^{d})\otimes\mathscr{D}_{\bullet\mathrm{b}}$.
@@
Intuitively, the logarithmic derivative $\mathsf{d}^{\nu}({x},\mathsf{s})$ represents the force received by the tagged particle at $x$ from (infinitely many) other particles $\mathsf{s}=\sum_{i}\delta_{s^{i}}$. Informally, $\mathsf{d}^{\nu}$ is defined as the differential of the logarithm of the Hamiltonian if the potential gives the equilibrium state $\nu$. One can easily calculate the logarithmic derivatives of canonical Gibbs measures with Ruelle’s class potentials using the DLR equation [25]. Even if the DLR equation does not hold, $\mathsf{d}^{\nu}({x},\mathsf{s})$ can be well defined—this is the case for the Coulomb potentials.
The logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ plays a vital role because, when it exists, it allows us to derive the stochastic dynamics of an infinite-particle system with the reversible equilibrium distribution $\nu$. This connection leads to the formulation of solutions through ISDEs, underscoring the power of this mathematical concept in understanding complex systems [25, 26, 37]:
$dX_{t}^{i}=dB_{t}^{i}+\frac{1}{2}\mathsf{d}^{\nu}(X_{t}^{i},X_{t}^{i\diamond})dt\quad(i\in\mathbb{N}),$ (2.11)
where $X_{t}^{i\diamond}=\sum_{j\neq i}^{\infty}\delta_{X_{t}^{j}}$.
In [25], it was proved that if $\nu$ has a logarithmic derivative $\mathsf{d}^{\nu}$ and the quasi-Gibbs property, then ISDE (2.11) has a weak solution and the associated unlabeled dynamics $\mathsf{X}_{t}=\sum\delta_{X_{t}^{i}}$ is a $\nu$-reversible diffusion. Here, the quasi-Gibbs property is a concept such that, loosely speaking, on each bounded domain, the regular conditional probability $\mu(\cdot|\mathsf{s}(S_{R})=m,\pi_{S_{R}^{c}}(\mathsf{s})=\pi_{S_{R}^{c}}(\mathsf{y}))$ of $\mu$ has a local density with upper and lower inequalities to positive functions presented by given potentials [26]. The quasi-Gibbs property is a generalization of the concept of the DLR equation and applies to Coulomb and Riesz potentials.
These two properties are satisfied by random point fields with logarithmic interaction potentials such as the Ginibre, sinβ, Airyβ, $\beta=1,2,4$, and Bessel random point fields [25, 26, 27, 35, 10]. The second author proved that, for any random point fields, only the existence of a logarithmic derivative is sufficient for constructing such dynamics [37]. The construction of stochastic dynamics using logarithmic derivatives is universal for reversible dynamics that arise from general random point fields with Ruelle’s class potentials and random matrix theory.
The existence of the logarithmic derivative does not guarantee the uniqueness of solutions for ISDEs, a matter that remains unresolved. The resulting stochastic dynamics cannot be fully understood without more detailed information about the logarithmic derivative. Therefore, obtaining a clear and explicit representation of the logarithmic derivative is essential.
If the logarithmic derivative has an explicit expression, we can prove that the solution to (2.11) is a pathwise unique, strong solution [34, 16]. This paper proves the existence of the logarithmic derivative for the Coulomb random point field $\mu$, obtains an explicit expression to establish the uniqueness of the strong solution, and investigates other properties of stochastic dynamics. Our method is novel and robust. It can apply to a broad class of random point fields with potentials beyond the Coulomb potentials.
2.2. Assumptions
We begin by introducing the assumptions used throughout the paper. For $\epsilon>0$, set $\mathbb{R}_{\epsilon}^{2d}=\{({x},{y})\in(\mathbb{R}^{d})^{2};|{x}-{y}|>\epsilon\}$.
i For each $N\in\mathbb{N}$, assume $\Phi^{N}\in C_{b}^{2}(O_{N})$, with $\Phi^{N}({x})=\infty$ for ${x}\notin\overline{O_{N}}$. Suppose $\Phi\in C^{2}(\mathbb{R}^{d})$ and
$\lim_{N\to\infty}\Phi^{N}({x})=\Phi({x})\quad\text{ in $C_{b}^{2}(S_{R})$ for each $R\in\mathbb{N}$},$ (2.12)
$\Delta\Phi({x})\geq 0\quad\text{ for all }{x}\in\mathbb{R}^{d}.$ (2.13)
ii Let $\Psi^{N}:(\mathbb{R}^{d})^{2}\to\mathbb{R}\cup\{\infty\}$ satisfy
$\Psi^{N}({x},{y})=0\text{ for }({x},{y})\notin\overline{O_{N}\times O_{N}},$ (2.14)
$\lim_{N\to\infty}$ $\Psi^{N}({x},{y})=\Psi({x}-{y})\text{ in $C_{b}^{2}(S_{Q}^{2}\cap\mathbb{R}_{\epsilon}^{2d})$ for each $Q\in\mathbb{N},\,\epsilon>0$},$ (2.15)
@@ 2.2. Assumptions
where $\delta_{{y}}$ is the delta measure at y{y}, $c_{2.1}^{N}$ is a positive constant, and $\psi^{N}\in C_{b}(O_{N}\times O_{N})$ satisfies, for some constant $c_{2.2}$,
[…]
((A2)) i $\mu^{N}$ has the labeled density $\mathfrak{m}^{N}$ with respect to $d\mathbf{{x}}^{N}=\prod_{i=1}^{N}d{x}^{i}$ defined by
$\mathfrak{m}^{N}(\mathbf{{x}}^{N})=\frac{1}{\mathscr{Z}^{N}}\exp\Big{(}-\beta\Big{\{}\sum_{i=1}^{N}\Phi^{N}({x}^{i})+\sum_{j,k=1,\,j\neq k}^{N}\Psi^{N}({x}^{j},{x}^{k})\Big{\}}\Big{)},$ (2.18)
where $0<\beta<\infty$ and $\mathscr{Z}^{N}$ is the normalizing constant:
$\mathscr{Z}^{N}=\frac{1}{N!}\int_{(\mathbb{R}^{d})^{N}}\exp\Big{(}-\beta\Big{\{}\sum_{i=1}^{N}\Phi^{N}({x}^{i})+\sum_{j,k=1,\,j\neq k}^{N}\Psi^{N}({x}^{j},{x}^{k})\Big{\}}\Big{)}d\mathbf{{x}}^{N}.$ (2.19)
ii There exists a subsequence $\{\mu^{N_{\mathfrak{n}}}\}$ converging weakly in $\mathsf{S}$ to some random point field $\mu$:
$\mu=\lim_{\mathfrak{n}\to\infty}\mu^{N_{\mathfrak{n}}}\quad\text{ weakly in }\mathsf{S}$ (2.20)
such that
$\limsup_{\mathfrak{n}\to\infty}\int_{\mathsf{S}}\mathsf{s}(S_{Q})\mu^{N_{\mathfrak{n}}}(d\mathsf{s})<\infty\quad\text{ for each }Q\in\mathbb{N},$ (2.21)
$\int_{\mathsf{S}}\mathsf{s}(S_{R})\mu(d\mathsf{s})=O(R^{\ell_{0}})\quad\text{ as }R\to\infty\text{ for some $\ell_{0}>0$},$ (2.22)
$\mu(\{\mathsf{s}(\mathbb{R}^{d})=\infty\})=1.$ (2.23)
Remark 2.3.
i From (2.21), $\{\mu^{N_{\mathfrak{n}}}\}_{\mathfrak{n}\in N_{\mathfrak{n}}}$ is tight, and
$\int_{\mathsf{S}}\mathsf{s}(S_{Q})\mu(d\mathsf{s})<\infty\quad\text{ for each }Q\in\mathbb{N}.$
ii Condition (2.22) requires that the expected number of particles in $S_{R}$ grows at most polynomially. This ensures that the logarithmic derivative of $\mu$ can be constructed. Assumption (2.23) simply excludes the trivial finite-particle case.
iii We do not assume a priori that $\mu$ admits a density or correlation functions; these follow from our assumptions (see Theorem 4.1).
[…]
Let $\overline{S}_{R}=\{s\in\mathbb{R}^{d};\lvert s\rvert\leq R\}$ and $S_{R+\epsilon}=\{s\in\mathbb{R}^{d};\lvert s\rvert<R+\epsilon\}$.
For a function $f$, we set $f^{(\mathbf{i})}=\partial^{k}f/\partial{x}^{\mathbf{i}}$ for $\mathbf{i}\in\mathbf{I}(k)$, where
@@
$\mathscr{X}_{R}(l)=\{{x}\in S_{R};\det[\,{x}_{\mathbf{j}}^{\mathbf{i}}\,]_{\mathbf{i},\mathbf{j}\in\mathbb{I}(l)}\not=0\},\quad{x}_{\mathbf{j}}^{\mathbf{i}}=({x}_{1}^{j_{1}})^{i_{1}}\cdots({x}_{d}^{j_{d}})^{i_{d}}.$ (2.25)
((A3)) For $\ell_{0}$ in (2.22), there exists an $\ell\in\mathbb{N}$ satisfying $\ell>\ell_{0}-d$ and the following:
i For each $R\in\mathbb{N}$ and $\epsilon>0$,
$\lim_{\mathfrak{n}\to\infty}\Psi^{N_{\mathfrak{n}}}({x},y)=\Psi({x}-{y})\quad\text{ in $C_{b}^{\ell+2}(\overline{S}_{R}^{2}\cap\mathbb{R}_{\epsilon}^{2d})$}.$ (2.26)
ii We set $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$. For each $R\in\mathbb{N}$ and $\epsilon>0$,
$\max_{\mathbf{i}\in\mathbf{I}(\ell+1)}\int_{\mathsf{S}}\sum_{{y}^{i}\notin S_{R+\epsilon}}^{\infty}\sup_{\mathfrak{n}\in\mathbb{N}}\big{\lVert}(-\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}({x},{y}^{i})\big{\rVert}_{C^{1}(\overline{S}_{R})}d\mu(\mathsf{y})<\infty.$ (2.27)
Here, we set $(-\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}({x},{y})=0$ for ${y}\notin O_{N_{\mathfrak{n}}}$ although $\Psi^{N_{\mathfrak{n}}}({x},{y})$ may be discontinuous in y{y} on $\partial O_{N_{\mathfrak{n}}}$.
iii For $\mu$-a.s. $\mathsf{y}=\sum_{i}\delta_{{y}^{i}}$ and each $R\in\mathbb{N}$,
$\lim_{\mathfrak{n}\to\infty}\sum_{{y}^{i}\not\in\overline{S}_{R}}^{N_{\mathfrak{n}}}\nabla_{{x}}\Psi^{N_{\mathfrak{n}}}({x},{y}^{i})-\nabla\Psi({x}-{y}^{i})=0\quad\text{ for some }{x}\in\mathscr{X}_{R}(\ell),$ (2.28)
$\lim_{\mathfrak{n}\to\infty}\max_{\mathbf{i}\in\mathbf{I}(\ell+1)}\sum_{{y}^{i}\not\in\overline{S}_{R}}^{N_{\mathfrak{n}}}\Big{\lVert}(\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}({x},{y}^{i})-(\nabla\Psi)^{(\mathbf{i})}({x}-{y}^{i})\Big{\rVert}_{C^{1}(\overline{S}_{R})}=0.$ (2.29)
@@
When $\Psi^{N}=\Psi$, ((A3)) automatically holds. Indeed, i and iii are obvious and $i$ follows from (1.1), (2.22), and $\ell>\ell_{0}-d$ (see Lemma 5.6).
These three assumptions are sufficient to obtain an explicit formula for the coefficient of the ISDE arising from $\mu$ and to construct the stochastic dynamics as the unique, strong solution of the ISDE.
To obtain the ISDE in translation-invariant form, we further assume the following.
((A4)) The following convergence holds in $L_{\mathrm{loc}}^{2}(\mu^{[1]})$, for $\mu^{[1]}$-a.e. $({x},\mathsf{s})$, or in $C^{1}(\overline{S}_{Q,\mathsf{s},\epsilon})$ for each $\mu$-a.s. $\mathsf{s}=\sum_{i}\delta_{s^{i}}$ and $\epsilon>0$:

### v4

@@ 2.1. Set up and logarithmic derivatives
i The logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ is an $\mathbb{R}^{d}$-valued function such that $\mathsf{d}^{\nu}\in L_{\mathrm{loc}}^{1}(\nu^{[1]})$ and that, for all $f\in\mathscr{D}_{\circ\mathrm{b}}^{[1]}$,
$\int_{\mathbb{R}^{d}\times\mathsf{S}}f(x,\mathsf{s})\mathsf{d}^{\nu}(x,\mathsf{s})\nu^{[1]}(dxd\mathsf{s})=-\int_{\mathbb{R}^{d}\times\mathsf{S}}\nabla_{{x}}f(x,\mathsf{s})\nu^{[1]}(dxd\mathsf{s}).$ (2.2)
ii The $\bullet$-logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ is the $\mathbb{R}^{d}$-valued function defined similarly to the logarithmic derivative, but with $\mathscr{D}_{\circ\mathrm{b}}^{[1]}$ replaced by $\mathscr{D}_{\bullet\mathrm{b}}^{[1]}$.
Intuitively, the logarithmic derivative $\mathsf{d}^{\nu}({x},\mathsf{s})$ represents the force received by the tagged particle at $x$ from (infinitely many) other particles $\mathsf{s}=\sum_{i}\delta_{s^{i}}$. Informally, $\mathsf{d}^{\nu}$ is defined as the differential of the logarithm of the Hamiltonian if the potential gives the equilibrium state $\nu$. One can easily calculate the logarithmic derivatives of canonical Gibbs measures with Ruelle’s class potentials using the DLR equation [27]. Even if the DLR equation does not hold, $\mathsf{d}^{\nu}$ can be well defined—this is the case for the Coulomb potentials.
The logarithmic derivative $\mathsf{d}^{\nu}$ of $\nu$ plays a vital role because, when it exists, it allows us to derive the stochastic dynamics of an infinite-particle system with the reversible equilibrium distribution $\nu$. This connection leads to the formulation of solutions $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$ through ISDEs, underscoring the power of this concept in understanding complex systems [27, 28, 36]:
$X_{t}^{i}-X_{0}^{i}=B_{t}^{i}+\frac{1}{2}\int_{0}^{t}\mathsf{d}^{\nu}(X_{u}^{i},\sum_{j\neq i}^{\infty}\delta_{X_{u}^{j}})du,\quad i\in\mathbb{N},$ (2.3)
where $\{B^{i}\}$ are independent standard Brownian motions.
2.2. Assumptions
We begin by introducing the assumptions used throughout the paper.
For $\epsilon>0$, we define $S_{R,\epsilon}^{2}=\{({x},{y})\in S_{R}^{2};|{x}-{y}|>\epsilon\}$.
i Suppose $\Phi\in C^{2}(\mathbb{R}^{d})$. For each $N,R\in\mathbb{N}$, let $\Phi^{N}\in C_{b}^{2}(O_{N})$ satisfy
$\Phi({x})$ $=\lim_{N\to\infty}\Phi^{N}({x}),$ $\text{in $C_{b}^{2}(S_{R})$},$ (2.4)
$\Phi^{N}({x})$ $=\infty,$ ${x}\notin O_{N},$ (2.5)
$\Delta\Phi^{N}({x})$ $\geq 0,$ ${x}\in O_{N}.$ (2.6)
ii Let $\Psi^{N}:(\mathbb{R}^{d})^{2}\to\mathbb{R}\cup\{\infty\}$ satisfy, for each $R\in\mathbb{N},\,\epsilon>0$,
$\Psi({x}-{y})$ $=\lim_{N\to\infty}\Psi^{N}({x},{y}),$ $\text{in $C_{b}^{2}(S_{R,\epsilon}^{2})$},$ (2.7)
$\Psi^{N}({x},{y})$ $=0,$ $({x},{y})\notin O_{N}^{2},$ (2.8)
[…]
where $c_{2.1}^{N}$ is a positive constant, $\delta_{{y}}$ is the delta measure at y{y}, and $\psi^{N}\in C_{b}(O_{N}^{2})$ satisfies, for some constant $c_{2.2}$,
[…]
((A2)) i The RPF $\mu^{N}$ admits the labeled density given in (1.3), where $\beta>0$ is arbitrary and fixed throughout this paper.
ii There exist an $\ell_{0}>0$ and a constant $c_{2.3}$ such that,
[…]
iii There exists a weakly convergent subsequence of $\{\mu^{N}\}$, denoted by $\{\mu^{N_{\mathfrak{n}}}\}$, with limit $\mu$ such that
$\mu\bigl(\{\mathsf{s}(\mathbb{R}^{d})=\infty\}\bigr)=1.$ (2.12)
@@ 2.2. Assumptions
i From (2.11), the family $\{\mu^{N}\}_{N\in\mathbb{N}}$ is tight, and
@@
This estimate is used in the construction of the logarithmic derivative of $\mu$.
ii Assumption (2.12) merely excludes the trivial finite-particle case.
iii We do not assume a priori that $\mu$ admits a density or correlation functions; these properties follow from the above assumptions (see Proposition 4.1).
For a function $f(x)$, we set $f^{(\mathbf{i})}=\partial^{k}f/\partial{x}^{\mathbf{i}}$ for $\mathbf{i}\in\mathbf{I}(k)$, where
@@
$\mathscr{S}_{R}^{l}=\{{x}\in S_{R};\det[\,{x}_{\mathbf{j}}^{\mathbf{i}}\,]_{\mathbf{i},\mathbf{j}\in\mathbb{I}(l)}\not=0\},\ {x}_{\mathbf{j}}^{\mathbf{i}}=({x}_{1}^{j_{1}})^{i_{1}}\cdots({x}_{d}^{j_{d}})^{i_{d}}.$ (2.15)
We introduce the following assumption.
When $\Psi^{N}=\Psi$, ((A3)) below holds automatically. Indeed, (2.16), (2.18) and (2.19) are obvious, and (2.17) follows from Lemma 5.8.
((A3)) For the subsequence $N_{\mathfrak{n}}$ in ((A2)) iii, there exists an $\ell\in\mathbb{N}$ satisfying $\ell>\ell_{0}-d$ such that the following hold:
i For each $\mathbf{i}\in\mathbf{I}(\ell+1)$, $R\in\mathbb{N}$, and $\epsilon>0$,
$\Psi({x}-{y})=\lim_{\mathfrak{n}\to\infty}\Psi^{N_{\mathfrak{n}}}({x},y)\quad\text{ in $C_{b}^{\ell+2}(S_{R,\epsilon}^{2})$},$ (2.16)
$\int_{\mathsf{S}}\Big(\sum_{s^{i}\not\in S_{R+\epsilon}}^{\infty}\sup_{\mathfrak{n}\in\mathbb{N}}\big\lVert(-\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}({x},s^{i})\big\rVert_{C^{1}(\overline{S}_{R}\cap O_{N_{\mathfrak{n}}})}\Big)\mu(d\mathsf{s})<\infty.$ (2.17)
ii For each $\mathbf{i}\in\mathbf{I}(\ell+1)$, $R\in\mathbb{N}$, and $\mu$-a.s. $\mathsf{s}$,
$\lim_{\mathfrak{n}\to\infty}\sum_{s^{i}\not\in\overline{S}_{R}}^{N_{\mathfrak{n}}}\Big\lVert(\nabla\Psi)^{(\mathbf{i})}({x}-s^{i})-(\nabla_{{x}}\Psi^{N_{\mathfrak{n}}})^{(\mathbf{i})}({x},s^{i})\Big\rVert_{C^{1}(\overline{S}_{R})}=0,$ (2.18)
$\lim_{\mathfrak{n}\to\infty}\sum_{s^{i}\not\in\overline{S}_{R}}^{N_{\mathfrak{n}}}\Big(\nabla\Psi({x}-s^{i})-\nabla_{{x}}\Psi^{N_{\mathfrak{n}}}({x},s^{i})\Big)=0\ \text{ for some }{x}\in\mathscr{S}_{R}^{\ell}.$ (2.19)
These three assumptions are sufficient to obtain an explicit formula for the coefficient of the ISDE arising from $\mu$ and to construct the stochastic dynamics as the unique, strong solution of the ISDE. To obtain the ISDE in translation-invariant form, we further assume the following:
((A4)) The following convergence holds in $L_{\mathrm{loc}}^{2}(\mu^{[1]})$, for $\mu^{[1]}$-a.e. $({x},\mathsf{s})$, or in $C^{1}(\overline{S}_{R,\mathsf{s},\epsilon})$ for $\mu$-a.s. $\mathsf{s}$ and $\epsilon>0$:
Here $\overline{S}_{R,\mathsf{s},\epsilon}=\{{x}\in\overline{S}_{R}\,;\,\lvert{x}-s^{i}\rvert\geq\epsilon\text{ for all }i\}$.
@@
The assumptions imposed on the limit RPF are mild. We believe that they allow the construction of many examples of RPFs with Coulomb interactions to which our theorems apply.
Unlabeled and labeling maps. Let $\mathfrak{u}\!:\!\big(\sqcup_{k=0}^{\infty}(\mathbb{R}^{d})^{k}\big)\sqcup(\mathbb{R}^{d})^{\mathbb{N}}\!\to\!\mathscr{M}(\mathbb{R}^{d})$ be the unlabeled map defined by
$\mathfrak{u}\big((x^{i})_{i}\big)=\sum_{i}\delta_{x^{i}},$ (2.21)
where $\mathscr{M}(\mathbb{R}^{d})$ is the set of all measures on $(\mathbb{R}^{d},\mathscr{B}(\mathbb{R}^{d}))$. If $k=0$, we set $\mathfrak{u}(x)=\mathsf{0}$ for $x\in(\mathbb{R}^{d})^{0}$ by convention, where $\mathsf{0}$ denotes the zero measure.
[…]
$\mathfrak{u}\circ\mathfrak{l}=\mathrm{id}.$
[…]
By (2.12), we have $\mathfrak{l}(\mathsf{s})\in(\mathbb{R}^{d})^{\mathbb{N}}$ for $\mu$-a.s. $\mathsf{s}$. For $\mathsf{s}\in\mathsf{S}$, we often write $\mathfrak{l}(\mathsf{s})=(s^{i})_{i}$. For $\mathsf{y}\in\mathsf{S}$, we define $\mathsf{y}^{N}=\mathsf{0}$ if $\mathsf{y}=\mathsf{0}$ and, for $\mathsf{y}\neq\mathsf{0}$,
$\mathsf{y}^{N}=\sum_{i=1}^{\mathsf{y}(\mathbb{R}^{d})\wedge N}\delta_{\mathfrak{l}^{i}(\mathsf{y})}.$ (2.23)
Conditional RPFs and couplings. Let $\overline{S}_{R}^{c}=\{\lvert s\rvert>R\}$ as before. Let $\mu_{R,\mathsf{y}^{N}}^{N}$ be the regular conditional probability of $\mu^{N}$ given by
$\mu_{R,\mathsf{y}^{N}}^{N}=\mu^{N}(\cdot\mid\pi_{\overline{S}_{R}^{c}}(\mathsf{s})=\pi_{\overline{S}_{R}^{c}}(\mathsf{y}^{N})).$ (2.24)
From ((A1)) and ((A2)), the labeled density $\mathfrak{m}^{N}$ is continuous on $O_{N}^{N}$. Hence, $\mu_{R,\mathsf{y}^{N}}^{N}$ is well defined for all $\mathsf{y}\in\mathsf{S}$. Moreover, $\mu_{R,\mathsf{y}^{N}}^{N}$ admits the $k$-density functions $\mathfrak{m}_{R,\mathsf{y}}^{N,k}$. For $\mu$-a.s. $\mathsf{y}$, the sequence $\{\mathsf{y}^{N}\}_{N\in\mathbb{N}}$ defines a coupling for the family of RPFs $\{\mu_{R,\mathsf{y}^{N}}^{N}\}_{N\in\mathbb{N}}$.
2.3. Explicit formulas for logarithmic derivatives: Theorem 2.1
In this subsection, we present explicit formulas for the logarithmic derivatives. This result is crucial for proving the uniqueness of solutions and the existence of strong solutions to the ISDE.
Let $\mathbf{I}(k)$ as in (2.14). Let $\ell$ be as in ((A3)). For $\mathbf{i}\in\mathbf{I}(\ell+1)$, let
$A_{\mathbf{i},\mathsf{s}}({x})$ $=\sum_{s^{i}\in\overline{S}_{R}^{c}}^{\infty}\int_{0}^{1}(-\nabla\Psi)^{(\mathbf{i})}(t{x}-s^{i})(\ell+1)(1-t)^{\ell}\,dt.$ (2.25)
Here we use the integral form of the Taylor expansion of $\nabla\Psi({x}-s^{i})$ around x=0{x}=0. By (2.17), the infinite sum in (2.25) converges absolutely in $C^{1}(\overline{S}_{R})$ for $\mu$-a.s. $\mathsf{s}$ (see Proposition 5.1). We define the function $\mathscr{R}_{R,\mathsf{s}}^{\ell}\in C^{1}(\overline{S}_{R})$ by
$\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})$ $=\sum_{\mathbf{i}\in\mathbf{I}(\ell+1)}\frac{{x}^{\mathbf{i}}}{i_{1}!\cdots i_{d}!}A_{\mathbf{i},\mathsf{s}}({x}).$
Let $\eta_{R,\mathsf{s}}^{\mathbf{i}}$ be the constant in (5.25). Let $\mathbb{I}(\ell)=\sqcup_{k=0}^{\ell}\mathbf{I}(k)$. Let
$\overline{S}_{Q,\epsilon}(\mathsf{s})=\{{x}\in\overline{S}_{Q}\,;\,\lvert{x}-s^{i}\rvert\geq\epsilon\text{ for all }i\},\quad\mathbb{R}^{d}_{\neq}(\mathsf{s})=\cap_{i}\{x\in\mathbb{R}^{d}\,;\,x\neq s^{i}\}$
Let $\mu^{[1]}$ be the one-reduced Campbell measure of $\mu$.
Theorem 2.1 (Explicit formulas for $\mathsf{d}^{\mu}$).
i Assume ((A1))((A3)). Then $\mu$ admits a $\bullet$-logarithmic derivative $\mathsf{d}^{\mu}\in L_{\mathrm{loc}}^{2}(\mu^{[1]})$, which is given by, for each $R\in\mathbb{N}$ and for $\mu^{[1]}$-a.e. $({x},\mathsf{s})\in S_{R}\times\mathsf{S}$,
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})+\sum_{s^{i}\in\overline{S}_{R}}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}+\sum_{\mathbf{i}\in\mathbb{I}(\ell)}\eta_{R,\mathsf{s}}^{\mathbf{i}}{x}^{\mathbf{i}}+\mathscr{R}_{R,\mathsf{s}}^{\ell}({x})\Big).$ (2.26)
ii Assume ((A1))((A3)). Then, for $\mu$-a.s. $\mathsf{s}$, each $Q\in\mathbb{N}$, $\mathbf{i}\in\mathbb{I}(\ell)$, and $\epsilon>0$,
$\lim_{R\to\infty}\eta_{R,\mathsf{s}}^{\mathbf{i}}=0,\quad\lim_{R\to\infty}\big\lVert\mathscr{R}_{R,\mathsf{s}}^{\ell}\big\rVert_{C^{1}(\overline{S}_{Q})}=0,$ (2.27)
$\mathsf{d}^{\mu}({x},\mathsf{s})=\beta\Big(-\nabla\Phi({x})+\lim_{R\to\infty}\sum_{s^{i}\in\overline{S}_{R}}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}\Big)\quad\text{in $C^{1}(\overline{S}_{Q,\epsilon}(\mathsf{s}))$}$ (2.28)
and $\mathsf{d}^{\mu}({x},\mathsf{s})$ admits a $\mu^{[1]}$-version that is locally Lipschitz in $x$ on $\mathbb{R}^{d}_{\neq}(\mathsf{s})$.
iii Assume ((A1))((A4)). Then $\mathsf{d}^{\mu}$ can be expressed as
$\mathsf{d}^{\mu}({x},\mathsf{s})=\lim_{R\to\infty}\beta\sum_{\lvert{x}-s^{i}\rvert\leq R}\frac{{x}-s^{i}}{\lvert{x}-s^{i}\rvert^{d}}.$ (2.29)
The convergence in (2.29) is understood in the sense specified by ((A4)).